Solve the equation. Beware of extraneous solutions. (15/x) + ((9x-7)/(x+2)) =9

Answer by stanbon(48535)

[15/x] + [(9x-7)/(x+2)]= [9/1]
LCM = x(x+2)
Multiply both sides by the LCM to get:
15(x+2)+(9x-7)x = 9x(x+2)
15x+30+9x^2-7x = 9x^2+18x
Simplify:
10x=30
x=3
Cheers,
Stan H.